The average atomic weight of oxygen (O) is 15.9994 amu. It has three naturally occurring isotopes, 16O, 17O and 18O, and 0.037 percent of oxygen is made up of 17O. If the atomic weights are 16O = 15.995 amu, 17O = 16.999 amu and 18O = 17.999 amu, what are the abundances of the other two isotopes?


1
Expert's answer
2021-01-20T03:15:06-0500

16.999 x 0.037/100 = 0,00628963 g.

15.9994 - 0,00628963 = 15,9931104 g.

15.995x + 17.999y= 15,9931104

x+y = 0,99963

x= 0,99963-y.

x= 0.002.

y= 0.99763